OpenStudy (tordtom):

How many molecules are in 50mL of water and how do you calculate how many atoms of hydrogen and oxygen are in it?

OpenStudy (mww):

Here you need to know the relationship between moles, mass and number of atoms/molecules. First number of moles can be expressed in two ways 1) as a mass \[n = \frac{ mass }{ molar ~mass } = \frac{ m }{ M }\] 2) as number of atoms/molecules described using Avogadro's number \[n = \frac{ number~ of~ particles }{ Avogadro's~ number } = \frac{ N }{ N_A } = \frac{ N }{ 6.022 \times 10^{23} }\] For 50ml of water, you need to convert this first to a mass. We can assume the density of the water is 1 g/cm^3 = 1g/mL, unless otherwise stated. Thus 50mL can be estimated as 50 g of water. Now we must convert this mass quantity into moles. \[n = \frac{ m }{ M } = \frac{ 50g }{ 18.016g/mol } = 2.775 ~moles\] Now that it is in moles we use the particle mole formula: \[n = \frac{ N }{ 6.022 \times 10^{23} } = 2.775 ~so~ N = 2.775 \times 6.022 \times 10^{23} = 1.7 \times 10^{24} \]

OpenStudy (mww):

The answer above tells you how many water molecules are in 50 g of water. To find the individual atoms, look at the chemical formula of water \[H_2O ~has~ 2H + O\] So each molecule of water corresponds to 2 atoms of hydrogen and one atom of oxygen. Another way to consider this If you had 5 boxes of fruit and each box had 2 apples and 1 banana, how much of each fruit do you have? 1 molecule = 1 x 2 apples + 1x 1 banana, so 5 molecules = 5 x 2 apples + 5 x 1 banana = 10 apples and 5 bananas.

OpenStudy (mww):

Remember molar units is the bridge between masses and particles so be able to use both mass and particle formulae by changing to moles. \[\frac{ m }{ M } = n = \frac{ N }{ N_A }\]